Construct 90%, 95%, and 99% confidence intervals to estimate μ from the following data. State the point estimate. Assume the data come from a normally distributed population.

12.1 11.6 11.9 12.3 12.5 11.4 12.0
11.7 11.8 12.1


Appendix A Statistical Tables


(Round the intermediate values to 4 decimal places. Round your answers to 2 decimal places.)

90% confidence interval:

μ


95% confidence interval:

μ


99% confidence interval:

μ



The point estimate is

.
Answer #1

solution & civen that ? point Estimate is . The mean oso - 120.5 2 12.05 10 storderd dewahon csson - ID:1845 = 0.43 D 90 % co23 95% confidence interval 2 20:05 X2= 0.025 difa (0-1)=9 +212= 2.26 = from table a coctai2. Sirm aulcst * 205/80 - (12-05-22Ponoalomy 19.11) = Toinhoo +50.7) >m > binno - 50 •31) =